# Trustworthy Transfer Learning: A Survey

### v1

@@ Abstract
Transfer learning aims to transfer knowledge or information from a source domain to a relevant target domain. In this paper, we understand transfer learning from the perspectives of knowledge transferability and trustworthiness. This involves two research questions: How is knowledge transferability quantitatively measured and enhanced across domains? Can we trust the transferred knowledge in the transfer learning process? To answer these questions, this paper provides a comprehensive review of trustworthy transfer learning from various aspects, including problem definitions, theoretical analysis, empirical algorithms, and real-world applications. Specifically, we summarize recent theories and algorithms for understanding knowledge transferability under (within-domain) IID and non-IID assumptions. In addition to knowledge transferability, we review the impact of trustworthiness on transfer learning, e.g., whether the transferred knowledge is adversarially robust or algorithmically fair, how to transfer the knowledge under privacy-preserving constraints, etc. Beyond discussing the current advancements, we highlight the open questions and future directions for understanding transfer learning in a reliable and trustworthy manner [^1] This work was mainly completed when Jun Wu was a PhD student at UIUC..
1 Introduction
@@
Standard machine learning assumes that training and testing samples are independently and identically drawn (IID). With this IID assumption, modern machine learning models (e.g., deep neural networks (?)) have achieved promising performance in a variety of high-impact applications. However, this IID assumption is often violated in real-world scenarios, especially when samples are collected from different sources and environments (?, ?). Transfer learning has been introduced to tackle the distribution shifts between training (source domain) and testing (target domain) data sets. In contrast to standard machine learning involving samples from a single domain, transfer learning focuses on modeling heterogeneous data collected from different domains. The intuition behind transfer learning is to bridge the gap between source and target data by discovering and transferring their shared knowledge (?). Compared to learning from the target domain alone, the transferred knowledge could significantly improve the prediction performance on the target domain, especially when the target domain has limited or no labeled data (?, ?). In recent decades, by instantiating the learning models with modern neural networks, a deep transfer learning paradigm has been introduced with enhanced transferability capabilities (?).
@@
As illustrated in (?), transfer learning is a general term to describe the transfer of knowledge or information from source to target domains. Depending on the data and model assumptions, it can lead to various specific problem settings, such as data-level knowledge transfer (domain adaptation (?, ?, ?), out-of-distribution generalization (?, ?), and self-taught learning (?)) with available source samples and model-level knowledge transfer (fine-tuning (?), source-free adaptation (?, ?), knowledge distillation (?)) with a pre-trained source hypothesis. The generalization performance of transfer learning techniques under various data and model assumptions has been studied over the past decades (?, ?, ?, ?). In addition to generalization performance, it is crucial to understand the trustworthiness (?) of the transferred knowledge in the transfer learning process, especially in safety-critical applications such as self-driving cars and medical diagnosis. It is explained (?) that “trust is the relationship between a trustor and a trustee: the trustor trusts the trustee”. In the context of transfer learning, the trustor can be the owners/users/regulators of either source or target domain. The trustee can be the transfer learning model itself, or the knowledge transferred from the source domain to the target domain. As summarized in earlier studies (?, ?, ?), various trustworthiness properties can encourage the “trustor” to trust the “trustee” in real scenarios, including adversarial robustness, privacy, fairness, transparency, etc. Therefore, in this paper, we focus on trustworthy transfer learning (?) that aims to understand transfer learning from the perspective of both knowledge transferability and knowledge trustworthiness.
Figure 1: A motivating example of trustworthy transfer learning
@@
Figure 1 provides a motivating example of trustworthy transfer learning in precision agriculture (?). In this example, a target farmer aims to train a model over the collected sorghum data. The task is to predict the biochemical traits (e.g., Nitrogen content, chlorophyll, etc.) of sorghum samples using the leaf hyperspectral reflectance (?, ?). Nevertheless, it is expensive and time-consuming to collect the labeled training samples. A feasible solution is to leverage knowledge from a relevant maize data set collected by a source farmer. This transfer learning process might involve several trustworthy concerns from source and target farmers. To name a few, will the privacy of source data be leaked in transfer learning? How does the poisoned and biased source knowledge negatively affect the prediction performance on the target domain? What is the fundamental trade-off between transfer performance and trustworthy properties? More generally, from the perspective of data and AI model markets (?), this emphasizes the importance of establishing trustworthiness between customers and sellers when purchasing AI models and sharing personal data.
@@
This survey provides a comprehensive review of state-of-the-art theoretical analysis and algorithms for trustworthy transfer learning. More specifically, we summarize recent theories and algorithms for understanding knowledge transferability from two aspects: IID and non-IID transferability. IID transferability assumes that the samples within each domain are independent and identically distributed. In this scenario, we review three major quantitative metrics for evaluating the transferability across domains, including (data-level) distribution discrepancy, (task-level) take diversity, and (model-level) transferability estimation. In contrast, non-IID transferability considers a more relaxed assumption that the samples within each domain can be interdependent, e.g., connected nodes in graphs (?), word occurrence in texts (?), temporal observations in time series (?), etc. We then review how transferability across domains can be quantitatively measured and enhanced in these complex scenarios. In addition to knowledge transferability, we also review the impact of trustworthiness on transfer learning techniques, including privacy, adversarial robustness, fairness, transparency, etc. Finally, we will highlight the open questions and future directions of trustworthy transfer learning.
@@ 2 Preliminaries
Figure 2: Overview of trustworthy transfer learning (best viewed in color)
2.1 Notation
@@
2.2 Problem Definition
Transfer learning (?) refers to the knowledge or information transfer from the source domain to the target domain such that the prediction performance on the target domain could be significantly improved as compared to learning from the target domain alone. Moreover, in the following definition, we generalize standard transfer learning (?) to trustworthy transfer learning (?).
Definition 1 (Trustworthy Transfer Learning)
@@ 2.2 Problem Definition
The source and target domains might involve different learning tasks (?, ?) or data modalities (?, ?). There are two key components in trustworthy transfer learning: knowledge transferability and trustworthiness. Specifically, knowledge transferability measures how the source knowledge can be successfully transferred to the target domain. In contrast, knowledge trustworthiness aims to answer whether transfer learning techniques provide reliable and trustworthy results in the target domain. Figure 2 provides a brief summarization of trustworthy transfer learning regarding knowledge transferability and trustworthiness (discussed in Section 3 and Section 4).
3 Knowledge Transferability
@@
3.1 IID Transferability
Here we summarize different transferability indicators, including distribution discrepancy, task diversity, and transferability estimator.
3.1.1 Distribution Discrepancy
Distribution discrepancy quantitatively measures the distribution shifts between two domains in the distribution space, when the source and target domains share the same input and output spaces (this scenario is also known as domain adaptation (?)). There are different types of distribution shifts (?), including covariate shift (?) (feedback covariate shift (?, ?)), label/target shift (?, ?), concept shift (?), etc. The covariate shift holds that the conditional probability $p(y|x)$ is shared across domains, but the marginal $p(x)$ is different. Label shift assumes that the conditional probability $p(x|y)$ is shared across domains, while the marginal label distribution $p(y)$ changes. Concept shift involves changes in the conditional probability $p(x|y)$ (or $p(y|x)$) changes, while the marginal distribution $p(y)$ (or $p(x)$) is fixed. The integral probability metric (IPM) (?, ?, ?) is a general framework for quantifying the difference between two distributions, and it can be instantiated by various statistical discrepancy measures (?), e.g., total variation distance, Wasserstein distance, maximum mean discrepancy (?), etc.
Definition 2 (Integral Probability Metric (?))
@@ 3.1.1 Distribution Discrepancy
[…]
@@
The concept of distribution discrepancy is the key to theoretically understanding how knowledge can be transferred from source to target domains. For example, the seminal work of ? (?) derives a generalization bound for domain adaptation using a tractable $\mathcal{H}\Delta\mathcal{H}$-divergence. Many follow-up works have developed refined generalization bounds by introducing various discrepancy measures. The following theorem provides a unified view of such generalization bounds based on a notion of discrepancy $d(\mathcal{D}_{S},\mathcal{D}_{T})$.
Theorem 1 (Unified Generalization Bound)
@@
where $\Omega$ represents the redundant terms (depending on how $d(\mathcal{D}_{S},\mathcal{D}_{T})$ is instantiated), e.g., the difference of labeling functions across domains (?, ?), the complexity of hypothesis space $\mathcal{H}$ (?, ?), number of training samples (?, ?), etc.
@@
We have the following observations regarding this unified view of generalization error. (1) The complexity of the class of hypothesis functions $\mathcal{H}$ plays a crucial role in deriving tight generalization error bounds. Various metrics have been applied to quantify this complexity (?), including the Vapnik-Chervonenkis (VC) dimension (?, ?, ?, ?), Rademacher complexities (?, ?, ?, ?, ?), covering number (?, ?), etc. We refer the reader to the survey (?) for more discussion. (2) Generally, the discrepancy $d(\mathcal{D}_{S},\mathcal{D}_{T})$ measures the difference between source and target distributions over the joint space $\mathcal{X}\times\mathcal{Y}$, when the distribution shifts occur across domains. In practice, it is commonly seen that the discrepancy $d(\mathcal{D}_{S},\mathcal{D}_{T})$ is defined over the input space $\mathcal{X}$ (when no label information is available in unsupervised domain adaptation), or over the joint space $\mathcal{X}\times\mathcal{Y}$ (when (pseudo)labels for target samples are available). The first type of $d(\mathcal{D}_{S},\mathcal{D}_{T})$, defined over the input space $\mathcal{X}$, is often associated with the covariate shift assumption (?) or the redundant term indicating the difference of labeling functions across domains, when deriving the generalization error bound. In the following, we summarize several commonly used discrepancy metrics $d(\mathcal{D}_{S},\mathcal{D}_{T})$.
@@
  • Total Variation Distance (?): The total variation distance (also referred to as $L^{1}$ divergence) between source and target domains can be defined as
    $d_{\mathrm{TV}}(\mathcal{D}_{S},\mathcal{D}_{T})=\sup_{B\in\mathcal{B}}\left|\mathrm{P}_{S}[B]-\mathrm{P}_{T}[B]\right|$ (2)
@@
    where $\mathcal{B}$ is the set of measurable subsets under $\mathrm{P}_{S}$ and $\mathrm{P}_{T}$. It is shown (?) that the total variation distance can be considered as a special case of the integral probability metric.
@@
  • $\mathcal{H}\Delta\mathcal{H}$-divergence (?, ?): It is illustrated (?, ?) that the empirical estimate of the total variation distance in Eq. (2) has two limitations. First, it cannot be accurately estimated from finite samples of arbitrary distributions in practice. Second, it results in loose generalization bounds due to involving a supremum over all measurable subsets. To address these limitations, ? (?) and ? (?) introduce the following $\mathcal{H}\Delta\mathcal{H}$-divergence:
    $d_{\mathcal{H}\Delta\mathcal{H}}(\mathcal{D}_{S},\mathcal{D}_{T})=\sup_{h,h^{\prime}\in\mathcal{H}}\left|\mathrm{P}_{S}\left[h(x)\neq h^{\prime}(x)\right]-\mathrm{P}_{T}\left[h(x)\neq h^{\prime}(x)\right]\right|$ (3)
@@
  • Discrepancy Distance (?): ? (?) extend the $\mathcal{H}\Delta\mathcal{H}$-divergence to a more general discrepancy distance for measuring distribution differences.
    $d_{\mathrm{disc}}(\mathcal{D}_{S},\mathcal{D}_{T})=\max_{h,h^{\prime}\in\mathcal{H}}\left|\mathbb{E}_{x\sim\mathrm{P}_{S}}\left[\ell\left(h(x),h^{\prime}(x)\right)\right]-\mathbb{E}_{x\sim\mathrm{P}_{T}}\left[\ell\left(h(x),h^{\prime}(x)\right)\right]\right|$ (4)
    where $\ell(\cdot,\cdot)$ denotes a general loss function (though the derived generalization bounds require that $\ell(\cdot,\cdot)$ is symmetric and obeys the triangle inequality). When using 0-1 classification loss, this discrepancy distance exactly recovers the $\mathcal{H}\Delta\mathcal{H}$-divergence. The discrepancy distance can be flexibly applied to compare distributions across various tasks, e.g., regression (?, ?).
  • $\mathcal{Y}$-discrepancy (?): It is notable that the discrepancy distance in Eq. (4) quantifies the difference between two marginal distributions over $\mathcal{X}$, when the ground-truth labeling function is unknown in the target domain. Later, ? (?) further extend the discrepancy distance to the $\mathcal{Y}$-discrepancy, which is defined over $\mathcal{X}\times\mathcal{Y}$ as follows.
    $d_{\mathcal{Y}}(\mathcal{D}_{S},\mathcal{D}_{T})=\sup_{h\in\mathcal{H}}\left|\mathbb{E}_{(x,y)\sim\mathrm{P}_{S}}\left[\ell\left(h(x),y\right)\right]-\mathbb{E}_{(x,y)\sim\mathrm{P}_{T}}\left[\ell\left(h(x),y\right)\right]\right|$ (5)
    In practice, this discrepancy can be estimated using pseudo labels of the target data, when there are no labeled data in the target domain (?, ?).
  • Margin Disparity Discrepancy (?): ? (?) extend the notion of discrepancy distance in Eq. (4) to margin disparity discrepancy (MDD) in multi-class classification settings. Specifically, MDD involves two key refinements: (1) the use of a margin-based loss function, and (2) the formulation of the discrepancy over both a hypothesis space $\mathcal{H}$ and a specific classifier $h$.
    $d_{\mathrm{MDD}}(\mathcal{D}_{S},\mathcal{D}_{T})=\sup_{h^{\prime}\in\mathcal{H}}\left|\mathbb{E}_{x\sim\mathrm{P}_{S}}\left[\Phi_{\rho}\left(\rho_{h^{\prime}}(x,h(x))\right)\right]-\mathbb{E}_{x\sim\mathrm{P}_{T}}\left[\Phi_{\rho}\left(\rho_{h^{\prime}}(x,h(x))\right)\right]\right|$ (6)
@@
  • $f$-divergence (?): Building on the margin disparity discrepancy (MDD) (?), ? (?) further develop a generic notion of the discrepancy based on the variational characterization of $f$-divergence (?). Specifically, the $f$-divergence is bounded by

### v2

@@ Abstract.
Transfer learning aims to transfer knowledge or information from a source domain to a relevant target domain. In this paper, we understand transfer learning from the perspectives of knowledge transferability and trustworthiness. This involves two research questions: How is knowledge transferability quantitatively measured and enhanced across domains? Can we trust the transferred knowledge in the transfer learning process? To answer these questions, this paper provides a comprehensive review of trustworthy transfer learning from various aspects, including problem definitions, theoretical analysis, empirical algorithms, and real-world applications. Specifically, we summarize recent theories and algorithms for understanding knowledge transferability under (within-domain) IID and non-IID assumptions. In addition to knowledge transferability, we review the impact of trustworthiness on transfer learning, e.g., whether the transferred knowledge is adversarially robust or algorithmically fair, how to transfer the knowledge under privacy-preserving constraints, etc. Beyond discussing the current advancements, we highlight the open questions and future directions for understanding transfer learning in a reliable and trustworthy manner.
copyright: ccdoi: 10.1613/jair.1.17602journalvolume: 84article: 20publicationmonth: 11journalyear: 2025
1. Introduction
@@
Standard machine learning assumes that training and testing samples are independently and identically drawn (IID). With this IID assumption, modern machine learning models (e.g., deep neural networks (lecun2015deep)) have achieved promising performance in a variety of high-impact applications. However, this IID assumption is often violated in real-world scenarios, especially when samples are collected from different sources and environments (TL_survey_2010; wu2024distributional). Transfer learning has been introduced to tackle the distribution shifts between training (source domain) and testing (target domain) data sets. In contrast to standard machine learning involving samples from a single domain, transfer learning focuses on modeling heterogeneous data collected from different domains. The intuition behind transfer learning is to bridge the gap between source and target data by discovering and transferring their shared knowledge (TL_survey_2010). Compared to learning from the target domain alone, the transferred knowledge could significantly improve the prediction performance on the target domain, especially when the target domain has limited or no labeled data (A_distance; task_diversity). In recent decades, by instantiating the learning models with modern neural networks, a deep transfer learning paradigm has been introduced with enhanced transferability capabilities (YosinskiCBL14).
@@
As illustrated in (TL_survey_2010), transfer learning is a general term to describe the transfer of knowledge or information from source to target domains. Depending on the data and model assumptions, it can lead to various specific problem settings, such as data-level knowledge transfer (domain adaptation (A_distance; DANN; discrepancy_distance), out-of-distribution generalization (BlanchardLS11; MuandetBS13), and self-taught learning (RainaBLPN07)) with available source samples and model-level knowledge transfer (fine-tuning (ShachafBG21), source-free adaptation (SHOT; AghbalouS23), knowledge distillation (hinton2015distilling)) with a pre-trained source hypothesis. The generalization performance of transfer learning techniques under various data and model assumptions has been studied over the past decades (task_diversity; 0002CZG19; minami2023transfer; MohriSS19). In addition to generalization performance, it is crucial to understand the trustworthiness (eshete2021making) of the transferred knowledge in the transfer learning process, especially in safety-critical applications such as autonomous driving and medical diagnosis. It is explained (Varshney2022) that “trust is the relationship between a trustor and a trustee: the trustor trusts the trustee”. In the context of transfer learning, the trustor can be the owners/users/regulators of either the source or the target domain. The trustee can be the transfer learning model itself, or the knowledge transferred from the source domain to the target domain. As summarized in earlier studies (eshete2021making; Varshney2022; kaurURD23), various trustworthiness properties can encourage the “trustor” to trust the “trustee” in real scenarios, including adversarial robustness, privacy, fairness, transparency, etc. Therefore, in this paper, we focus on trustworthy transfer learning (wu2023trustworthy) that aims to understand transfer learning from the perspective of both knowledge transferability and knowledge trustworthiness.
Figure 1. A motivating example of trustworthy transfer learning
@@
Fig. 1 provides a motivating example of trustworthy transfer learning in precision agriculture (adve2024aifarms). In this example, a target farmer aims to train a model over the collected sorghum data. The task is to predict the biochemical traits (e.g., Nitrogen content, chlorophyll, etc.) of sorghum samples using the leaf hyperspectral reflectance (wang2023airborne; DINO). Nevertheless, it is expensive and time-consuming to collect the labeled training samples. A feasible solution is to leverage knowledge from a relevant maize data set collected by a source farmer. This transfer learning process might involve several trustworthy concerns from the source and target farmers. To name a few, will the privacy of source data be leaked in transfer learning? How does the poisoned and biased source knowledge negatively affect the prediction performance on the target domain? What is the fundamental trade-off between transfer performance and trustworthy properties? More generally, from the perspective of data and AI model markets (pei2023data), this emphasizes the importance of establishing trustworthiness between customers and sellers when purchasing AI models and sharing personal data.
@@
This survey provides a comprehensive review of state-of-the-art theoretical analysis and algorithms for trustworthy transfer learning. More specifically, we summarize recent theories and algorithms for understanding knowledge transferability from two aspects: IID and non-IID transferability. IID transferability assumes that the samples within each domain are independent and identically distributed. In this scenario, we review three major quantitative metrics for evaluating the transferability across domains, including (data-level) distribution discrepancy, (task-level) take diversity, and (model-level) transferability estimation. In contrast, non-IID transferability considers a more relaxed assumption that the samples within each domain can be interdependent, e.g., connected nodes in graphs (GCN), word occurrence in texts (LeeDS18), temporal observations in time series (VRADA), etc. We then review how transferability across domains can be quantitatively measured and enhanced in these complex scenarios. In addition to knowledge transferability, we also review the impact of trustworthiness on transfer learning techniques, including privacy, adversarial robustness, fairness, transparency, etc. Finally, we will highlight the open questions and future directions of trustworthy transfer learning.
@@ 2. Preliminaries
Figure 2. Overview of trustworthy transfer learning (best viewed in color)
2.1. Notation
@@
2.2. Problem Definition
Transfer learning (TL_survey_2010) refers to the knowledge or information transfer from the source domain to the target domain such that the prediction performance on the target domain could be significantly improved as compared to learning from the target domain alone. Moreover, in the following definition, we generalize standard transfer learning (TL_survey_2010) to trustworthy transfer learning (wu2023trustworthy).
Definition 2.1 (Trustworthy Transfer Learning).
@@ 2.2. Problem Definition
The source and target domains might involve different learning tasks (TL_survey_2010; task_diversity) or data modalities (ShenLDSKNT23; IGLUE). There are two key components in trustworthy transfer learning: knowledge transferability and trustworthiness. Specifically, knowledge transferability measures how the source knowledge can be successfully transferred to the target domain. In contrast, knowledge trustworthiness aims to answer whether transfer learning techniques provide reliable and trustworthy results in the target domain. Fig. 2 provides a brief summarization of trustworthy transfer learning regarding knowledge transferability and trustworthiness (discussed in Section 3 and Section 4).
3. Knowledge Transferability
@@
3.1. IID Transferability
Here, we summarize different transferability indicators, including distribution discrepancy, task diversity, and transferability measures.
3.1.1. Distribution Discrepancy
Distribution discrepancy quantitatively measures the distribution shifts between two domains in the distribution space, when the source and target domains share the same input and output spaces (this scenario is also known as domain adaptation (TL_survey_2010)). There are different types of distribution shifts (wiles2022a), including covariate shift (shimodaira2000improving) (feedback covariate shift (fannjiang2022conformal; prinster2023jaws)), label/target shift (lipton2018detecting; zhang2013domain), concept shift (redko2019advances), etc. The covariate shift holds that the conditional probability $p(y|x)$ is shared across domains, but the marginal $p(x)$ is different. Label shift assumes that the conditional probability $p(x|y)$ is shared across domains, while the marginal label distribution $p(y)$ changes. Concept shift involves changes in the conditional probability $p(x|y)$ (or $p(y|x)$), while the marginal distribution $p(y)$ (or $p(x)$) is fixed. The integral probability metric (IPM) (muller1997integral; sriperumbudur2010hilbert; zhang2012generalization) is a general framework for quantifying the difference between two distributions, and it can be instantiated by various statistical discrepancy measures (sriperumbudur2010hilbert), e.g., total variation distance, Wasserstein distance, maximum mean discrepancy (MMD), etc.
Definition 3.1 (Integral Probability Metric (muller1997integral)).
@@ 3.1.1. Distribution Discrepancy
(1) $d_{\mathrm{IPM}}(\mathcal{D}_{S},\mathcal{D}_{T})=\sup_{h\in\mathcal{H}}\left|\int_{M}h~d\mathrm{P}_{S}-\int_{M}h~d\mathrm{P}_{T}\right|$
@@
The concept of distribution discrepancy is the key to theoretically understanding how knowledge can be transferred from source to target domains. For example, the seminal work of (A_distance) derives a generalization bound for domain adaptation using a tractable $\mathcal{H}\Delta\mathcal{H}$-divergence. Many follow-up works have developed refined generalization bounds by introducing various discrepancy measures. The following theorem provides a unified view of such generalization bounds based on a notion of discrepancy $d(\mathcal{D}_{S},\mathcal{D}_{T})$.
Theorem 3.2 (Unified Generalization Bound).
@@
where $\Omega$ represents the redundant terms (depending on how $d(\mathcal{D}_{S},\mathcal{D}_{T})$ is instantiated), e.g., the difference of labeling functions across domains (A_distance; f_divergence), the complexity of hypothesis space $\mathcal{H}$ (discrepancy_distance; zhang2012generalization), number of training samples (A_distance; redko2017theoretical), etc.
@@
We have the following observations regarding this unified view of generalization error. (1) The complexity of the class of hypothesis functions $\mathcal{H}$ plays a crucial role in deriving tight generalization error bounds. Various metrics have been applied to quantify this complexity (redko2019advances), including the Vapnik-Chervonenkis (VC) dimension (ben2006analysis; blitzer2007learning; A_distance; peng2019moment), Rademacher complexities (f_divergence; SCA; MDD; discrepancy_distance; mohri2012new), covering number (MDD; zhang2012generalization), etc. We refer the reader to the survey (redko2019advances) for more discussion. (2) Generally, the discrepancy $d(\mathcal{D}_{S},\mathcal{D}_{T})$ measures the difference between source and target distributions over the joint space $\mathcal{X}\times\mathcal{Y}$, when the distribution shifts occur across domains. In practice, it is commonly seen that the discrepancy $d(\mathcal{D}_{S},\mathcal{D}_{T})$ is defined over the input space $\mathcal{X}$ (when no label information is available in unsupervised domain adaptation), or over the joint space $\mathcal{X}\times\mathcal{Y}$ (when (pseudo)labels for target samples are available). The first type of $d(\mathcal{D}_{S},\mathcal{D}_{T})$, defined over the input space $\mathcal{X}$, is often associated with the covariate shift assumption (shimodaira2000improving) or the redundant term indicating the difference of labeling functions across domains, when deriving the generalization error bound. In the following, we summarize several commonly used discrepancy metrics $d(\mathcal{D}_{S},\mathcal{D}_{T})$.
@@
  • Total Variation Distance: The total variation distance (also referred to as $L^{1}$ divergence) (ben2006analysis) between source and target domains can be defined as
    (2) $d_{\mathrm{TV}}(\mathcal{D}_{S},\mathcal{D}_{T})=\sup_{B\in\mathcal{B}}\left|\mathrm{P}_{S}[B]-\mathrm{P}_{T}[B]\right|$
@@
    where $\mathcal{B}$ is the set of measurable subsets under $\mathrm{P}_{S}$ and $\mathrm{P}_{T}$. It is shown (sriperumbudur2010hilbert) that the total variation distance can be considered as a special case of the integral probability metric.
@@
  • $\mathcal{H}\Delta\mathcal{H}$-divergence: It is illustrated (ben2006analysis; A_distance) that the empirical estimate of the total variation distance in Eq. (2) has two limitations. First, it cannot be accurately estimated from finite samples of arbitrary distributions in practice. Second, it results in loose generalization bounds due to involving a supremum over all measurable subsets. To address these limitations, (ben2006analysis; blitzer2007learning) introduce the following $\mathcal{H}\Delta\mathcal{H}$-divergence:
    (3) $d_{\mathcal{H}\Delta\mathcal{H}}(\mathcal{D}_{S},\mathcal{D}_{T})=\sup_{h,h^{\prime}\in\mathcal{H}}\left|\mathrm{P}_{S}\left[h(x)\neq h^{\prime}(x)\right]-\mathrm{P}_{T}\left[h(x)\neq h^{\prime}(x)\right]\right|$
@@
  • Discrepancy Distance: (discrepancy_distance) extends the $\mathcal{H}\Delta\mathcal{H}$-divergence to a more general discrepancy distance for measuring distribution differences.
    (4) $d_{\mathrm{disc}}(\mathcal{D}_{S},\mathcal{D}_{T})=\max_{h,h^{\prime}\in\mathcal{H}}\left|\mathbb{E}_{x\sim\mathrm{P}_{S}}\left[\ell\left(h(x),h^{\prime}(x)\right)\right]-\mathbb{E}_{x\sim\mathrm{P}_{T}}\left[\ell\left(h(x),h^{\prime}(x)\right)\right]\right|$
    where $\ell(\cdot,\cdot)$ denotes a general loss function (though the derived generalization bounds require that $\ell(\cdot,\cdot)$ is symmetric and obeys the triangle inequality). When using 0-1 classification loss, this discrepancy distance exactly recovers the $\mathcal{H}\Delta\mathcal{H}$-divergence. The discrepancy distance can be flexibly applied to compare distributions across various tasks, e.g., regression (discrepancy_distance; cortes2011domain).
  • $\mathcal{Y}$-discrepancy: It is notable that the discrepancy distance in Eq. (4) quantifies the difference between two marginal distributions over $\mathcal{X}$, when the ground-truth labeling function is unknown in the target domain. Later, (mohri2012new) further extend the discrepancy distance to the $\mathcal{Y}$-discrepancy, which is defined over $\mathcal{X}\times\mathcal{Y}$ as follows.
    (5) $d_{\mathcal{Y}}(\mathcal{D}_{S},\mathcal{D}_{T})=\sup_{h\in\mathcal{H}}\left|\mathbb{E}_{(x,y)\sim\mathrm{P}_{S}}\left[\ell\left(h(x),y\right)\right]-\mathbb{E}_{(x,y)\sim\mathrm{P}_{T}}\left[\ell\left(h(x),y\right)\right]\right|$
    In practice, this discrepancy can be estimated using pseudo labels of the target data, when there are no labeled data in the target domain (long2018conditional; courty2017joint).
  • Margin Disparity Discrepancy: (MDD) extend the notion of discrepancy distance in Eq. (4) to margin disparity discrepancy (MDD) in multi-class classification settings. Specifically, MDD involves two key refinements: (1) the use of a margin-based loss function, and (2) the formulation of the discrepancy over both a hypothesis space $\mathcal{H}$ and a specific classifier $h$.
    (6) $d_{\mathrm{MDD}}(\mathcal{D}_{S},\mathcal{D}_{T})=\sup_{h^{\prime}\in\mathcal{H}}\left|\mathbb{E}_{x\sim\mathrm{P}_{S}}\left[\Phi_{\rho}\left(\rho_{h^{\prime}}(x,h(x))\right)\right]-\mathbb{E}_{x\sim\mathrm{P}_{T}}\left[\Phi_{\rho}\left(\rho_{h^{\prime}}(x,h(x))\right)\right]\right|$
@@
  • $f$-divergence: Building on the margin disparity discrepancy (MDD) (MDD), (f_divergence) further develop a generic notion of the discrepancy based on the variational characterization of $f$-divergence (nguyen2010estimating). Specifically, the $f$-divergence is bounded by
    $d_{f}(\mathcal{D}_{S},\mathcal{D}_{T})=\int p_{T}(x)\phi\left(\frac{p_{S}(x)}{p_{T}(x)}\right)dx\geq\sup_{T\in\mathcal{T}}\mathbb{E}_{x\sim\mathrm{P}_{S}}\left[T\left(x\right)\right]-\mathbb{E}_{x\sim\mathrm{P}_{T}}\left[\phi^{*}\left(T\left(x\right)\right)\right]$
    where $\phi(\cdot)$ is a convex lower semi-continuous function that satisfies $\phi(1)=0$. $\phi^{*}$ is the conjugate function of $\phi$. $\mathcal{T}$ is a set of measurable functions. Based on this observation,  (f_divergence) define a notion of $f$-divergence guided discrepancy as follows.
    (7) $d_{\phi}(\mathcal{D}_{S},\mathcal{D}_{T})=\sup_{h^{\prime}\in\mathcal{H}}\left|\mathbb{E}_{x\sim\mathrm{P}_{S}}\left[\ell\left(h(x),h^{\prime}(x)\right)\right]-\mathbb{E}_{x\sim\mathrm{P}_{T}}\left[\phi^{*}\left(\ell\left(h(x),h^{\prime}(x)\right)\right)\right]\right|$
  • Generalized Discrepancy: In contrast, (generalized_discrepancy; cortes2019adaptation) generalize the discrepancy distance in Eq. (4) using reweighting techniques. That is, the difference between two distributions can be adjusted by multiplying the loss for each training example by a non-negative weight (huang2006correcting; cortes2008sample; zhang2013domain). Formally, for any hypothesis-dependent reweighting function $U_{h}$, the generalized discrepancy is defined as follows.
    (8) $d_{\mathrm{DISC}}(\mathcal{D}_{S},\mathcal{D}_{T})=\max_{h\in\mathcal{H},h^{\prime\prime}\in\mathcal{H}^{\prime\prime}}\left|\mathbb{E}_{x\sim\hat{\mathrm{P}}_{S}}\left[{U}_{h}(x)\cdot\ell\left(h\left(x\right),f_{S}\left(x\right)\right)\right]-\mathbb{E}_{x\sim\hat{\mathrm{P}}_{T}}\left[\ell\left(h(x),h^{\prime\prime}(x)\right)\right]\right|$
    where $\hat{\mathrm{P}}_{S},\hat{\mathrm{P}}_{T}$ denote the empirical distributions of source and target domains, respectively, and $f_{S}$ denotes the source labeling function.
  • Rényi Divergence: Furthermore, (MansourMR09; cortes2010learning) derive the generalization bounds for adaptation approaches based on importance reweighting (e.g., sample reweighting for single-source adaptation (cortes2010learning) and domain reweighting for multi-source adaptation (MansourMR09)) based on the following Rényi divergence (renyi1961measures).
    (9) $d_{\alpha}(\mathcal{D}_{T}||\mathcal{D}_{S})=\frac{1}{\alpha-1}\log\sum_{x\in\mathcal{X}}\mathrm{P}_{T}(x)\left(\frac{\mathrm{P}_{T}(x)}{\mathrm{P}_{S}(x)}\right)^{\alpha-1}$
    where $\alpha\geq 0$.
  • Wasserstein Distance (Wasserstein): In general, fro any $p\geq 1$, the $p$-Wasserstein distance between two distributions can be defined as follows.
    (10) $d_{W_{p}}(\mathcal{D}_{S},\mathcal{D}_{T})=\left(\inf_{\pi\in\Pi(\mathrm{P}_{S},\mathrm{P}_{T})}\int c(x,x^{\prime})^{p}d\pi(x,x^{\prime})\right)^{1/p}$